### v1

@@ 3.4.1.1 The good
Inferring changes outside of the source and target APIs The migration in Figure 4 adds the async_mode argument to the SocketIO() function. This is interesting because the API is from the Flask-SocketIO library, which is not the source (eventlet) or target (gevent) libraries. However, the SocketIO function works with several libraries, eventlet being the default flasksocketio . This means, if 4o did not explicitly set the async_mode argument to gevent, the code would break after the migration.
@@
3.4.1.2 The Bad
We also analyze cases where the LLMs did not correctly migrate the code and present notable observations.
Failing to handle argument transformation Figure 5 shows an example where the source library retrying libretrying expects the time arguments in milliseconds, while the target library tenacity libtenacity expects them in seconds. The dataset includes 28 instances of this transformation, and all three LLMs failed to handle it in most cases, except for 4o, which handled three instances correctly, while the other two models correctly handled one instance each. This suggests that while the LLMs can map parameters, differences in expected format or units may not be properly managed.
Figure 5: Mini filing in argument transformation.
Replacing exception type We find cases where the LLM replaces generic exception types with specific ones, which can potentially fail to catch the same exceptions as the original code. This occurs in both migration-related and non-migration-related changes. For example, the LLM replaces an original Exception with ValueError. While using more specific exceptions is considered good practice cacho2014does , it does change the behavior of the code. Since modern development tools are often equipped to suggest better exception handling, it may be safer for an LLM-based migration tool to avoid changing existing exception types during a migration.
3.4.1.3 The ugly
We also observe the following problematic LLM behavior.
Extra changes LLMs are known for not always following the instructions provided in the prompt ye2023cognitive . Despite our prompt explicitly asking the LLMs to avoid any refactoring or non-migration changes, we find cases where an LLM, particularly LLama, completely removes parts of the code that do not include any source APIs. We speculate that this behavior may stem from the LLM attempting to optimize output size due to token limitations. Silently deleting code can be dangerous for the overall functionality of the system, and led to test failures in RQ2 for many cases.
Similarly, the LLMs sometimes refactor code that is not related to the migration, adopting a different style. While these improvements can be beneficial, they are unrelated to the migration task at hand. Future work could explore the use of an interactive IDE plugin that confirms the code changes before applying them, potentially allowing the developer to directly edit the changes.
While less “dangerous”, we also find that the LLMs remove code comments and docstrings from the source code, even though the comments remain relevant after migration.
Large file handling We observe that large files pose a challenge to LLama, given its smaller context size. Additionally, we observe that even the other two models often migrate an initial part of a large file but then incorrectly state that the remaining part does not require any changes.
3.4.2 Differences in evaluation setup
While the scope of RQ-A2 is limited due to the availability of the tests, we do find cases where the migrations are correct in RQ-A1 but incorrect in RQ-A2, or vice versa. To better understand these discrepancies, we manually inspect migrations with different results and provide some examples.
3.4.2.1 Library version compatibility
When evaluating RQ-A1, we mark a code change to be correct if it is correct as of the latest version of the library. However, in RQ-A2, we install the version the developers specified, and if it is not available, we resolve the library based on the migration commit date (see Section 3.3). Since we do not ask for a specific version number in the prompt, we notice that the LLMs usually pick APIs from the latest version of the target library. In some cases, this version discrepancy led to runtime errors when running tests in LLM’s code, which we report as incorrect in RQ-A2. Note that we consider version numbers moving forward when designing PyMigTool in Section 4.
3.4.2.2 Function parameter shadowing import name
ipaddr to ipaddress, google/capirca/eb768ea7/capirca/lib/nacaddr.py
Figure 6: Example of function parameter shadowing import name.
Notice in Figure 6 that the function parameter in the function IP on line 6 (Filepre) is named ipaddress, the same as the target library. The developer changed the parameter name to ip in Filedev. However, while Mini correctly replaces function IPNetwork with ip_network, it retained the original parameter name, causing it to shadow the new import name and leading to test failures. 4o changed the parameter name to ipaddress_str to avoid the shadowing issue.
3.4.2.3 Developer doing non-refactoring changes
Recall that in RQ-A1, we find cases where the developer makes non-refactoring changes in the same commit as the migration while the LLM only does the migration it is expected to do. We count this as a correct migration in RQ-A1. However, given the changed behavior (which is often reflected in the tests), the tests fail in RQ-A2.
Overall, the discrepancies show that both evaluation techniques provide different perspectives on the migration correctness.
3.4.3 Understanding the cost of LLM-based migration
The 314 migrations we use in our experiments involve a total of 258,198 lines of code, or 258.0 KLOC. The 10 runs per model makes it 2582.0 KLOC per model. The total cost of running our experiments was approximately \$15 for GPT-4o mini and \$420 for GPT-4o, which equates to \$0.01 and \$0.16 per KLOC, respectively. OpenAI has since reduced prices, so current costs would be lower. LLama 3.1 is free and open-source, though we used a pro subscription of Hugging Face (US\$9/month [^3] https://huggingface.co/pricing) to avoid any installation issues. While library migration is an infrequent task in a project’s lifetime, it requires significant developer effort when it happens. The cost of a developer’s time can easily exceed the cost of using an LLM, even GPT-4o, which can justify the cost of LLM-based tooling.
3.5 Threats to validity
Moha: I tried merging the two threats sections. But it because repetations of ”in phase one” and ”in phase two”. So I kept them separate for now.
3.5.1 Internal validity
Our evaluation involves two manual steps: validating correctness of code changes and labeling a non-migration related change as refactoring or non-refactoring. Manual activities are inherently subjective and may contain errors. We minimize this by having two authors independently perform the manual activities and resolving disagreements through discussion.
3.5.2 Construct validity
In RQ-A1, we analyze the code change categories based on Changedev, not on Changellm. The categories can differ if the LLM produces a different code change than the developer. Categorizing the LLM changes would require additional manual effort following the our previous work in PyMigTax pymigtax , which is beyond the scope of this work. That said, our observation is that the code/API differences in most cases are minor, and would not yield in a different PyMigTax category.
Tests passing may not always guarantee migration correctness (e.g., the tests do not cover certain edge cases). We mitigate this, to the extent possible, by selecting repositories with high test coverage in LABEL:rq:unseen and ensuring that the tests cover the source API usages.
3.5.3 External validity
We use three off the shelf LLMs. While using an LLM specifically tuned for the library migration task could potentially yield better results, the goal of this initial empirical study is not to develop a migration tool, but to understand the challenges and opportunities of using LLMs for library migration. Our results may be viewed as a lower bound of the performance of potentially more specialized models.
The libraries included in PyMigBench are all old-enough libraries where the LLMs are likely familiar with their APIs and API usage. This familiarity with APIs and API usage is precisely our hope when using LLMs for migration. However, being familiar with the APIs alone does not always guarantee being able to map them to each other or apply the right transformations on all code bases, which is why our evaluation is necessary. For newer libraries, additional information, e.g., API documentation, may be needed to guide the LLM in the migration process. This can be an interesting avenue for future work.
4 Phase two: PyMigTool: a tool for end-to-end Python library migration
In Section 3, we showed that LLMs can perform library migrations. However, we see two major challenges that make solely relying on LLMs in a practical tool ineffective: (1) We provide the LLM with only code files that directly use the source library. However, for a given project, there may be other files that depend on the migrated code and also require changes during the migration. (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration. In this phase, we build and evaluate PyMigTool, a CLI tool that automates the full library migration process for a full project by combining the power of LLMs with pre and post processing to address the above challenges.
PyMigTool first prepares the project that needs migration by setting up a virtual environment. It then runs an LLM to migrate the code, and finally performs post processing to improve the migration results. We run an extensive evaluation of PyMigTool on a set of 717 new, simulated migrations. Overall, PyMigTool is able to successfully migrate 32% of the migrations without any manual intervention. For a sample of the remaining not-fully-successful migrations, we find that a median of only 14% of the code lines in the project require manual changes to complete the migration, while the rest of the code is correctly migrated by PyMigTool.
4.1 Building PyMigTool
@@
PyMigTool takes as input the following: (1) the path to a python project where the migration should happen, (2) the source library name, (3) the target library name and version, (4) the python version in which the code should run, (5) paths to one or more requirements files, (6) and an LLM to use for the migration. A key criteria for the input project is that it contains tests with high coverage, as PyMigTool uses the tests to identify the source library usage as well as to verify the migration. While the migration will run regardless of the test coverage, migration results without test validation may not be as reliable.
PyMigTool works in three main steps: Preparation round, LLM migration round, and Post-processing round. The preparation round sets up a virtual environment using the requirements files, and runs the tests to get the baseline results. The LLM migration round migrates the project to use the target library instead of the source library using the LLM the user specified. Finally, if the LLM migration is not fully successful, PyMigTool performs some non-LLM post processing steps to improve the migration results.
4.1.1 Preparation round
@@
In the preparation round, PyMigTool first creates a virtual environment for the provided python version using the command python -m venv .venv, and then installs the required libraries using the command pip install -r <requirements-file-path> for each of the requirements file. It then installs the target library at the provided version using the pip install command. PyMigTool requires some additional packages, such as pytest, pytest-cov, which it also installs during this step.
@@
Once the environment is set up, PyMigTool runs the tests using the command python -m pytest on the project and saves the test report. We refer to this version of the test report as TestReportpremig, which is used as the baseline for the migration.
4.1.2 LLM migration round
In our preliminary exploration of LLMs (Section 3), we use PyMigBench where the files that require migration were already known. However, in a real-world scenario for a full project, we do not know which files require migration. Therefore, PyMigTool first identifies the files that require migration, and then uses the LLM to migrate these files.
4.1.2.1 Identifying files that require migration
PyMigTool migrates one whole file at a time, therefore, it is sufficient to identify which files use the source library, without necessarily identifying the exact source library usage within these files. PyMigTool combines two techniques to identify the files that require migration: it uses static analysis to identify import statements and uses dynamic analysis to generate a run-time-generated call graph of the project using profiling data.
geolib (library code)
1class Square:
2 def __init__(self, l):
3 self.l = l
4 self.name = "Square"
5
6 def area(self):
7 return self.l ** 2
serialize.py (uses library w/o import)
@@ 4.1.2.1 Identifying files that require migration
main.py (uses library with import)
@@
2from geo import Square
@@
serialize_test.py (test code)
1from serialize import to_dict
2from geo import Square
3
4def test_to_dict():
5 square = Square(4)
6 result = to_dict(square)
7 assert result == {"name": "Square", "area": 16}
Figure 7: Example of using a library with/without importing
Detecting import statements: If a file in the project imports a library, PyMigTool considers that the file uses the library, and thus require migration. In Python, a library may be imported using a name different than the library name. For example, the library PyYAML can be installed using the command pip install pyyaml, but it has to be imported as import yaml. The former, pyyaml, is the library name, and the latter, yaml, is the import name. Furthermore, a library can have more than one import names. PyMigTool only takes the source library name as input, therefore, it needs to find the import names of the source library in the project. To do this, PyMigTool uses the Johnnydep tool libjohnnydep , which extracts the import names from the wheel file of the library.
Next, PyMigTool uses the ast module to parse each Python file in the project and identify the import statements. It then compares the imported modules with the import names of the source library to identify the source library usage in the project.
In the example in Figure 7, the file main.py directly imports the library geolib, thus PyMigTool will identify this file as requiring migration. However, the file serialize.py does not import the library, therefore, will not be identified as requiring migration by the static analysis.
Detecting source library calls: In Figure 7, the file serialize.py does not import the library geolib, but both the main.py and serialize_test.py pass the geolib.Square object to the serialize.to_dict() function. That means when the code runs, the serialize.py file will use the geolib library, even though it does not import it. This shows that, identifying the import statements alone is not sufficient to identify all of the source library usage in the project. Accordingly, PyMigTool uses a run-time call graph to identify further source library usage in the project.
PyMigTool runs the tests in the project with profiling using the pprofile package libpprofile . The profiling data includes all the calls made during the test execution, and generates a report in the CallGrind format doccallgrind . A call information includes the caller function and the callee function, as well as the line number where the call happened. A function is identified by the path of the file in which the function is defined, and the qualified name which is unique within the file. PyMigTool uses this data to build a call graph of the project.
To know if a file requires migration, PyMigTool needs to determine whether any part of the file directly calls a source library function. A function can be defined in one of the three places: in python built-in modules, in the project, or in a third-party library. A Python system module file path is enclosed within angle brackets in the profile report; for example, <frozen os> indicates that the function is defined in the built-in os module. If a function is defined in a file under the project path (which PyMigTool takes as input), it considers this function to be in the project. PyMigTool also knows the path of the virtual environment where the third-party libraries are installed. Python virtual environments have a specific directory structure that PyMigTool leverages. For example, in Windows systems, the import name yaml would be in the <venv_path>/Lib/site-packages/yaml directory. Therefore, if a call is to a function that is defined in a file in this directory, PyMigTool can conclude that the function belongs to the yaml import name, i.e., the pyyaml library. In the call graph, we store the owner of each function, which can be project, system, or the import name of the library.
Given this constructed call graph, PyMigTool iterates over each call. If the caller is owned by the project and the callee is owned by the source library, then the file that contains the caller function requires migration and PyMigTool adds it to the list of files that will be passed to the LLM.
@@
Overall, in Figure 7, PyMigTool will identify the file main.py requiring migration using static analysis, and serialize.py requiring migration using dynamic analysis.
4.1.2.2 Migrating using an LLM
@@
PyMigTool can be configured to use any LLM that adheres to OpenAI’s APIs openaimodels , which has become a standard API for most models. For each file that requires migration, PyMigTool generates a prompt using the template in 2 and sends it to the LLM. The prompt is based on the prompt used in Section 4 (1), but enhanced with more specific instructions, and also includes the versions of the source and target libraries.
Listing 2: Prompt template used by PyMigTool
@@
1The following Python code currently uses the library <source-lib> version<source-ver>.
2Migrate this code to use the library <target-lib> version <target-ver> instead.
3
4**Instructions:**
51. **Explain the Changes**: Begin the output with a brief explanation of the specific changes you made to migrate from "<source-lib>" to "<target-lib>".
@@
62. **Provide the Modified Code**: After the explanation, present the modified code. Provide the entire code after migration even if only a part of it is changed.
7
8**Important Guidelines**:
@@
9- Only make changes directly related to migrating between "<source-lib>" and "<target-lib>".
10- Do not refactor, reformat, optimize, or alter the original coding style.
11- The code given to you is part of a larger application. Do not change the names of classes, functions, or variables, because it can break the application.
12
13Original code:
14<premig-code>
@@
The LLM response contains the migrated code, which PyMigTool extracts and saves. When the LLM finishes migrating all files, PyMigTool replaces the original files in the project with the migrated files. It then runs the tests and generates a test report, which we refer to as TestReportllmmig.
4.1.3 Post-processing round
PyMigTool performs two post-processing steps we describe below.
4.1.3.1 Merge skipped code
During the the preliminary experiments (Section 3), we found that it is common for the LLMs to skip some parts of the code in the response, especially if those parts do not require migration, despite being instructed to return the full migrated code (see paragraph 3.4.1.3). In this post-processing step, PyMigTool identifies such skipped code, and merges it with the migrated code.
@@ 4.1.3.1 Merge skipped code
For each file that was migrated in the LLM migration round, PyMigTool generates a diff between the original file and the migrated file using python’s difflib module. Then it attempts to detect which of the chunks were entirely removed from the original file. We make the following assumption to identify skipped code: 1. A skipped code must be mostly removal of lines. Because if the hunk has similar amount of additions and removals, it is likely that the hunk is replacing some code with other code, likely performing migration. 2. The skipped code should not use source library APIs, because this makes the hunk migration-related otherwise. 3. A skipped code should not be too small. This assumption avoids incorrectly considering small migration-related removals as skipped code.
We find the below specific baseline values for the above assumptions by trial-and-error to find a good balance between incorrectly merging migration related removals and incorrectly not merging skipped code:
  • Small (0-9 lines of removals): this hunk is not considered skipped irrespective of the number of additions.
  • Medium (10-19 lines of removals): if this hunk has at least one addition of line or use any source library APIs, it is not considered skipped.
  • Large (20 or more lines of removals): if this hunk has less than 90% of removed lines compared to total added+removed lines, it is not considered skipped.
The hunks contain the line numbers in the previous and new version of the file. Using the line numbers, PyMigTool inserts the removed lines in the migrated file, starting with the last removed hunk and going to the first removed hunk. The reverse order is important because inserting the lines would otherwise change the line numbers of the hunks that are inserted later. After inserting the skipped code, PyMigTool runs the tests again and generates a new test report, which we refer to as TestReportmerge.
4.1.3.2 Infer async transformation
In Python, a call to an asynchronous function must be awaited, i.e., preceded by the await keyword. Again, if a function call is awaited, the caller function must also be made asynchronous, i.e., it must be preceded by the async keyword.
When there is a migration from a synchronous library to an asynchronous library, some functions in the migrated code may become asynchronous, which requires their transitive calls to be awaited, and transitive callers to be made asynchronous as well. In this step, PyMigTool identifies the functions that were made asynchronous by the LLM (asynced functions), then find all other functions that needs to be made asynchronous (to-async functions) and all calls that need to be made await (to-await calls), and applies the changes to the project. Additionally, PyMigTool also decorates the corresponding test functions with the pytest.mark.asyncio decorator, and installs the pytest-asyncio package, which are required to correctly run asynchronous tests. Below we describe the steps in detail using the example in Figure 8.

### v2

@@ 3.4 Strengths of LLM for library migration
Different API styles
We find that the LLMs are able to handle transformation between different API styles. For example, the library argparse provides API functions for parsing command line arguments, while click provides decorators. Figure 4 shows a migration example between these libraries. Despite the completely different API styles, the three LLMs correctly migrate 94% of code changes between this library pair.
Before migration (using argparse)
4def parse_commands():
5 p = ArgumentParser()
6 p.add_argument("src", type=Path)
7 p.add_argument("dst", type=Path)
8 p.add_argument("usedir",
9 action = "store_true")
10 return parser.parse_args()
11
12def main():
13 args = parse_commands()
14 convert(args.src, args.dst, args.usedir)
4o’s migration (using click)
4@argument("src",
5 type=click.Path())
6@argument("dst",
7 type=click.Path())
8@argument("usedir", type=bool)
9def main(src, dst, usedir):
10 convert(src, dst, usedir)
11
12
13
14ÂX
Figure 4: 4o correctly migrating between libraries that use different API styles.
Inferring changes outside of the source and target APIs
We find that LLMs can infer some additional changes beyond the immediate APIs of the source or target library. For example, the migration in Figure 5 adds the async_mode argument to the SocketIO() function. This is interesting because the API is from the Flask-SocketIO library, which is not the source (eventlet) or target (gevent) libraries. However, the SocketIO function works with several libraries, eventlet being the default Flask (2024). This means, if 4o did not explicitly set the async_mode argument to gevent, the code would break after the migration.
Before migration (using eventlet)
3socketio = SocketIO(app)
4bootstrap = Bootstrap(app)
4o’s migration (using gevent)
3socketio = SocketIO(app, async_mode="gevent")
4bootstrap = Bootstrap(app)
@@
3.5 Challenges of using LLMs for library migration
From the results of the empirical study in Phase 1 and the above strengths, we see that LLMs have strong potential for library migration. However, we also noted several challenges that would need to be addressed if LLMs are to be used for practical tooling for library migration.
  1. Challenge-1
    Files requiring migration may not be known upfront: In our experiment, we use PyMigBench to identify the files that require migration that we will send to the LLM. However, in a real-world scenario, the files requiring migration may not be known upfront. Therefore, a tool needs to first identify the files that require migration before sending them to the LLM for migration.
  2. Challenge-2
    LLMs do not always return complete code: A common issue we observe across all three LLMs is that they often do not return the complete code. Instead, they skips parts of the code, usually commenting that the next block of code remains unchanged. This is especially when the skipped part does not require migration, despite our prompt explicitly asking them to return the entire code. While this is not necessarily incorrect, as the skipped code may not require any changes, the resulting code is incomplete and cannot be provided directly to the developer.
  3. Challenge-3
    Dependents of migrated files may also require changes: In our experiment, we provide the LLMs only the files that directly use the source library, as marked in PyMigBench. However, other files that depend on the migrated files may also require changes. This is especially true when functions are made asynchronous during the migration, as all parts of the code that use those functions must also be made asynchronous.
  4. Challenge-4
    Library version compatibility: We do not mention any library version requirements in the prompt (1), and we notice that the LLMs usually pick the APIs from the latest or recent versions of the target library. In few cases, this caused runtime errors when running tests in LLM’s code, which we report as incorrect in RQ-1.2. This is because the application or the Python version used in the project may not be compatible with the latest version of the target library. In an actual migration scenario, the developers may want to migrate to a specific version of the target library that is compatible with their project.
4 Phase 2: Building MigrateLib
In Section 3, we show that LLMs can perform library migrations. However, we also identify several challenges when using LLMs for library migration, especially when building an end-to-end tool that developers can use. In this section, we describe Phase 2 of our work, where we build MigrateLib while addressing the identified challenges through pre- or post-processing steps as follows:
  1. 1.
    Before migration, MigrateLib identifies the files that use the source library, and thus require migration to address Challenge-1.
  2. 2.
    MigrateLib allows the user to specify the source and target library version, and include this in the prompt, to address Challenge-4.
  3. 3.
    After migration, MigrateLib automatically adds back any skipped code to address Challenge-2.
  4. 4.
    After migration, MigrateLib identifies the functions that become asynchronous due to the migration, and finds all other functions in the projects that use these functions, and makes them asynchronous as well to address Challenge-3.
@@ 4 Phase 2: Building MigrateLib
MigrateLib works as a Command Line Interface (CLI) and takes the following as input: (1) the path to a Python project that needs library migration, (2) the source library name (the version is inferred), (3) the target library name and version, (4) the Python version in which the code should run, (5) paths to one or more requirements files, (6) and an LLM to use for the migration.
MigrateLib performs the migration in three main steps. First, it sets up a virtual environment for the project and identifies all files that require migration (Preparation step). In the next step, it prompts an LLM to migrate the identified files (LLM migration step). Finally, in the last step, it performs two post-processing steps to address Challenge-2 and Challenge-3 mentioned above to improve the migration results (Post-processing step).
We note that a key criteria for the input project is that it contains tests with high coverage, as MigrateLib uses the tests to identify the source library usage as well as to verify the migration. While the migration will run regardless of the test coverage, migration results without test validation may not be as reliable.
4.1 Preparation step
The preperation step handles three main preparation tasks needed before the actual migration:
Task 1: Set up environment
@@ 4.1 Preparation step
MigrateLib creates a virtual environment for the provided python version using the command python -m venv .venv, and then installs the required libraries using the pip install -r <requirements-file-path> command for each of the requirements file. It then installs the target library at the provided version using the pip install command. MigrateLib requires some additional packages, such as pytest, pytest-cov, which it also installs during this step. After the installation, MigrateLib derives the source library version by querying the installed package using the pip show <source-library-name> command.
Task 2: Run tests to get baseline
@@
Once the environment is set up, MigrateLib runs the tests using the command python -m pytest on the project and saves the test report. We refer to this version of the test report as TestReportpremig, which is used as the baseline for the migration.
Task 3: Identify files that require migration
This task addresses Challenge-1 from Section 3, where we need to identify the files that require migration before sending them to the LLM. One obvious way a file can use a library is by importing it. For example, in Figure 6, the file main.py imports the geolib library, thus it requires migration when migrating from geolib to another library. To identify such files, MigrateLib reads and parses all Python files in the project and looks for import statements of the source library. Note that a library can be imported using a different name than the library name; a library may also have multiple import names. Therefore, MigrateLib first finds all import names of the source library using the Johnnydep library Jeantine-Glenn (2025), which extracts the import names from the wheel file of the library.
A file may also use a library API without directly importing it. An example is shown in Figure 6. Notice that the to_dict() function in the serialize.py takes as input a shape object from the geolib library, even though it does not import the library. It uses the Shape.name attribute and Shape.area() method, which are part of the geolib library. To identify such files, MigrateLib builds a run-time call graph of the project by running the tests with profiling. It then analyzes the call graph to find all files that call any function from the source library.
serialize.py
@@
6 X
main.py
@@
2from geolib import Square
@@
Figure 6: Example of using a library without importing
@@
Overall, in Figure 6, MigrateLib will identify the file main.py requiring migration using static analysis, and serialize.py requiring migration using dynamic analysis.
4.2 LLM migration step
@@
MigrateLib can be configured to use any LLM that adheres to OpenAI’s APIs OpenAI (2024), which has become a standard API for most models. For each file that requires migration, MigrateLib generates a prompt using the template in 2 and sends the prompt to the LLM. The prompt is based on our original prompt used in Section 4 (1), but enhanced with more specific instructions. Especially, we include the versions of the source and target libraries to address Challenge-4 from Section 3.
Listing 2: Prompt template used by MigrateLib
@@
The following Python code currently uses the library <source-lib> version
<source-ver>. Migrate this code to use the library <target-lib> version
<target-ver> instead.
**Instructions:**
1. **Explain the Changes**: Begin the output with a brief explanation of the
specific changes you made to migrate from "<source-lib>" to "<target-lib>".
2. **Provide the Modified Code**: After the explanation, present the modified
@@
code. Provide the entire code after migration even if only a part of it is
changed.
**Important Guidelines**:
@@
- Only make changes directly related to migrating between "<source-lib>" and
"<target-lib>".
- Do not refactor, reformat, optimize, or alter the original coding style.
- The code given to you is part of a larger application. Do not change the names
of classes, functions, or variables, because it can break the application.
Original code:
<premig-code>
@@
The LLM response contains the migrated code, which MigrateLib extracts and saves. When the LLM finishes migrating all files, MigrateLib replaces the original files in the project with the migrated files. It then runs the tests and generates a test report, which we refer to as TestReportllmmig.
4.3 Post-processing Step 1: Merge skipped code
@@
We design this post-processing step to address Challenge-2 from Section 3. The goal is to add back any unchanged code that LLM skipped returning in its response. For each file that was migrated in the LLM migration step, MigrateLib generates a diff between the original file and the migrated file using Python’s difflib module. Then it attempts to detect which of the hunks were entirely removed from the original file. We make the following assumption to identify a hunk of skipped code:
  1. 1.
    The hunk should not use source library APIs such that it is not a migration-related change. Therefore, MigrateLib checks if any of the removed lines in the hunk uses any source library API. If yes, the hunk is not considered skipped.
  2. 2.
    The hunk must mostly include removal of lines. Our reasoning is that if the hunk has a similar number of additions and removals, it is likely that the hunk is replacing some code with other code, likely performing migration. If a hunk has more than 90% of removed lines compared to total added+removed lines, we consider it mostly removal.
  3. 3.
    The hunk should not be too small. This assumption avoids incorrectly considering small migration-related removals as skipped code. We consider a hunk as too small if it has less than 9 lines of code.
We determine the above criteria thresholds through experimentation on a set of migration examples. If a hunk is identified as skipped code by the above criteria, MigrateLib adds back the removed lines from the original file to the migrated file. After finishing the merging of skipped code, MigrateLib runs the tests again and generates a new test report, which we refer to as TestReportmerge.
4.4 Post-processing Step 2: Infer async transformation
This post-processing step addresses Challenge-3 from Section 3 by identifying functions that become asynchronous due to the migration, and making their transitive callers asynchronous as well. Below we describe the steps in detail using the example in Figure 7.